a flash unit for a camera has a capacitance of 1200μF.
what resistance is needed in this RC circuit if the flash isto charge to 90% of its full charge in 20s? R=
Answer #1
I think the answer should be as follow:
Given, C = 1200μF, t = 20s, R = ?
Using this equation, q(t) = Cε (1- e^-t/RC), set upthe charge to be 90% (0.90)
                        0.90Cε = Cε(1- e^-t/RC), now you can ignore the Cε since they're on each side
                       0.90 = (1- e^-t/RC)
                       e^-t/RC= 1- 0.90
                        e^-t/RC=0.10
                       ln e^-t/RC= ln 0.10
                       - t/RC = ln 0.10
       -20/((R)(1200μF)) =-2.3 , now solve for R
                  
                   R = 7246Ω
I hope that's what you're looking for.